OpenStudy (anonymous):

Multiply and simplify. show ur work (8x² )( 2x^5 )

OpenStudy (campbell_st):

the rule for multiplying the same base is to add the powers \[x^a \times x^b = x^{a + b}\] so in your question you have \[8x^2 \times 2x^5 = 8 \times 2 \times x^2 \times x^5\] add the powers and multiply the numbers to get the simplified version

OpenStudy (jhannybean):

beat meeeeeeeeeee hehehe